OpenStudy (anonymous):

solve for x: x^2=x-35

OpenStudy (anonymous):

Use the quadratic formula: \(x^2-x+35=0\) \(a=1,b=1,c=35\)

OpenStudy (amorfide):

b=-1 correct?

mathslover (mathslover):

right amorfide ... i think onmi did mistake in typing ..

OpenStudy (anonymous):

Yes, I did, lol. \(a=1, b=-1, c=35\)

OpenStudy (amorfide):

original equation: x²=x-35 re arrange the equation to get all terms onto one side and make it equal to zero you do this by subtracting and x on both sides and adding 35 onto both sides ( you do the opposite to the sign of the term/constant) the new equation given is:x²-x+35=0 the value of each thing in the equation can be given a letter in order so a,b,c a=1 b=-1 c=35 take note... x² and -x have a 1 and -1 infront of it... it still has a value anyway now that we have 3 values we can sub them into this formula